# CATP-LLM: Empowering Large Language Models for Cost-Aware Tool Planning

### v2

@@ 5.3 Ablation Study
Method Task Performance Execution Prices QoP %. of Valid Plans
w/o context augmentation 0.6538 0.0920 0.2458 100%
w/o intermediate feedback 0.7075 0.1643 0.2088 100%
w/o adaptive masking 0.6340 0.0855 0.2416 70.18%
CATP-LLM 0.6611 0.0794 0.2605 100%
Importance of different techniques in CAORL. The proposed fine-tuning algorithm CAORL plays an significant role in enabling CATP-LLM for cost-aware tool planning. Hence, we explore the contributions of various important techniques in CAORL to CATP-LLM, including context augmentation, intermediate feedback signals, and adaptive masking. As shown in Table 3, the context augmentation and intermediate feedback are essential for reducing execution costs, while adaptive masking ensures plan validness and improves generation effectiveness.
@@
Impacts of LLM parameter sizes. Finally, we explore the impacts of LLM sizes on CATP-LLM’s efficacy. We use TinyLlama-1B [52], Qwen-3B [17], and Llama2-7B [41] as backbone LLMs. As shown in Figure 7, Qwen-3B demonstrates comparable performance to Llama2-7B in sequential planning, but Llama2-7B significantly outperforms others in complex non-sequential tasks. This suggests that for relatively simple sequential tasks, smaller LLMs like Qwen-3B are sufficient for planning, while larger LLMs like Llama2-7B are recommended for more complex tasks. Due to GPU resource limits, we will leave the exploration of larger LLMs for future work.
@@ 6 Conclusion
In this paper, we propose the first cost-aware tool planning framework CATP-LLM. It features a tool planning language to enhance the LLM to create non-sequential plans of multiple branches for efficient concurrent execution. Building on this language, it further introduces a cost-aware offline reinforcement learning algorithm to fine-tune the LLM to optimize the trade-off between plan performance and costs. We then present OpenCATP, the first platform for cost-aware planning evaluation. Experiments on OpenCATP show that CATP-LLM significantly outperforms GPT-3.5/GPT-4 in cost-aware tool planning, even with a small 7B LLM as its backbone.
@@ A Details of CAORL
The details of the decision transformer [6] based learning algorithm in CAORL are described as follows. Formally, the LLM takes the historical return, state, and action sequences to predict the next action:
@@ B Details of OpenCATP
B.1 Evaluation Dataset in OpenCATP
OpenCATP features a range of compositional tasks for the comprehensive evaluation of cost-aware tool planning, which can be mainly categorized into sequential tasks and non-sequential tasks. For sequential tasks that can be solved with simple sequential plans, OpenCATP reuses the tasks originally provided in OpenAGI [10]. More importantly, considering that OpenAGI lacks challenging non-sequential tasks that must be solved with non-sequential plans, OpenCATP further constructs 24 additional non-sequential tasks with the method described in §$\S$4. Each sequential or non-sequential task is accompanied by 100 input data samples and corresponding ground-truth outputs. During experiments, we randomly select 40 samples of various input data sizes for each task. Table B.1 provides several concrete examples in OpenCATP.
Task Input Sample Ground Truth
@@ B.1 Evaluation Dataset in OpenCATP
Given a noisy blurry grayscale image, how to return the regular image step by step? [Uncaptioned image] [Uncaptioned image]
Given a noisy blurry grayscale image, how to return the caption in English step by step? [Uncaptioned image] A woman stands in the dining area at the table.
@@
Given a blurry noisy image, how to (1) return the regular image, and (2) return the label of the image in German step by step? [Uncaptioned image] (1) [Uncaptioned image]; (2) Schaf.
Given a blurry image, how to (1) return the label of the image in English, and (2) return the caption of the image in German step by step? [Uncaptioned image] (1) Braunbär; (2) Ein kräftiger Grizzly Bär ist im Hintergrund mit Gras zu sehen.
@@
B.2 Evaluation Metrics in OpenCATP
Execution prices. OpenCATP utilizes execution prices to comprehensively reflect the overall costs of tool plans, inspired by the Function as a Service (FaaS) platforms. FaaS is a category of cloud computing that charges consumers in an event-driven manner, where consumers only need to pay for the time and resources used to run their functions, with no prices paid when the functions are idle [36]. This “pay-as-you-go” model can result in significant savings, especially for applications with intermittent usage or short-lived tasks, contributing to the market success of FaaS. According to the recent statistics [18], the global market size of FaaS is estimated to be 17.70 billion USD in 2024 and is expected to reach 44.71 billion USD by 2029.
@@ B.2 Evaluation Metrics in OpenCATP
Here, $time$ denotes the execution time in milliseconds of tool $t$. $cpu_{cons}$ ($gpu_{cons}$) denotes the constant CPU (GPU) memory consumption in MB required for loading the tool. $cpu_{inst}$ ($gpu_{inst}$) denotes the instant CPU (GPU) memory consumption in MB caused by the computation of the tool. $price\_cpu_{cons}$, $price\_cpu_{inst}$, $price\_gpu_{cons}$, $price\_gpu_{inst}$ calculate the prices in USD associated with the respective memory consumptions. $price\_per\_run$ denotes the prices in USD charged for each execution of the tool. Table B.2 lists the settings of the pricing parameters [^3] Note that the existing FaaS platforms do not provide pricing strategy for GPU resources. We then set the GPU prices as three times of CPU prices, according to the article in https://news.rice.edu/news/2021/rice-intel-optimize-ai-training-commodity-hardware.
@@
Execution time. For sequential plan, the execution time is simply the sum of execution time of each tool in the plan. For example, for plan ① in Figure 1, the execution time is $0.18+3.46+0.13=3.77$(s). For non-sequential plan that can be executed in parallel, the execution time is defined as the maximum of the sum of tool execution time in each branch of the plan. For example, for plan ④ in Figure 1, the execution time is $\max(0.18+0.29+0.16,0.18+0.09)=\max(0.63,0.27)=0.63$ (s).
Task performance. In OpenCATP, we mainly use ViT Score [10] and BERT Score [53] to calculate the performance scores of a tool plan on the target task. Specifically, ViT Score is used for tasks that involve image outputs, which measures the cosine similarity between the image generated by the plan and the ground-truth image. BERT Score is used for tasks that involve text outputs, which measures the cosine similarity between the generated texts and ground-truth texts.
@@ C Details of Experiments
C.1 Implementation of CATP-LLM
@@
The details of implementing CATP-LLM on OpenCATP platform in the experiments are described as follows. For TPL, we assign each tool in OpenCATP with a unique tool token and a dependency token. The dependency token of a tool means accepting the outputs provided by this tool as inputs, as depicted in Figure 3. We also introduce a special dependency token $\langle task\rangle$ indicating accepting the task data as inputs. Note that in our experiments, we focus on resource dependencies, i.e. input-output dependencies. However, the concepts of TPL can extend to other types of dependencies, such as order dependencies where there exists strict execution orders between tools.
@@
As for CAORL, we use the execution prices and task performance defined in §$\S$§B.2 to calculate the execution costs and plan performance in the reward function. Besides, to implement the context augmentation scheme, we need to derive the cost attributes of each tool. To achieve this, we categorize the input data in OpenCATP into $k=4$ size levels using k-means clustering, with the optimal number of clusters determined by the elbow method. We then profile the execution time and CPU/GPU memory consumption of each tool across varying input sizes, following the method outlined in §$\S$3.2.1. Based on these statistics, we calculate the overall costs for each tool using the execution prices defined in Eq.(B.2), which ultimately yield the tool cost attributes.
C.2 Implementation of Baselines
@@
Zero-shot learning. For zero-shot learning, we design two types of prompts for for sequential planning and non-sequential planning, respectively. For sequential planning, we prompt the LLM to produce a tool sequence that can be sequentially executed to solve the target tasks. As for non-sequential planning, we instruct the LLM to generate a tool plan following the format similar to our TPL. This is because we find the LLM achieves poor performance in non-sequential planning without an appropriate format to generate non-sequential plans. Table C.1 illustrates the two types of prompts designed for zero-shot learning. Note that we also incorporate the average tool cost information into the prompts to instruct the LLM to create tool plans of low costs. The average tool cost information is derived from offline profiling results described in §$\S$§C.1.
@@
Few-shot learning. The few-shot learning share similar prompts with zero-shot learning, except that we handcraft several high-quality demonstrations as in-context examples to augment the LLM for plan generation. Table C.2 illustrates the prompts designed for few-shot learning.
@@
Note that the engines of GPT-3.5 and GPT-4 used for zero-shot learning, few-shot learning and HuggingGPT are gpt-3.5-turbo-0125 and gpt-4-turbo, respectively.
Scenarios Prompts
Sequential Planning (Command) You need to act as a policy model that, given a task and a set of tools, determines the sequence of tools that can be executed sequentially to solve the task. Your goal is to optimize the task performance while minimizing the execution costs.
(Tool Information) The information of each tool is provided as follows: Object Detection: This tool identifies the names of objects in an image. It is generally used for object identification in the iamges. The input and output types of this tool are image and text, respectively. It can accept inputs from tools “Image Super Resolution”, “Colorization”, “Image Deblurring”, or “Image Denoising”. Costs of Object Detection: On average, this tool takes about 175.73 milliseconds to run. It consumes approximately 352.19 MB of CPU memory and 449.37 MB of GPU memory. Image Deblurring: This tool can enhance the clarity of blurry images. It can be used for tasks that require improving image quality. Both the input and output types are images. It can accept inputs from tools “Image Super Resolution”, “Colorization”, or “Image Denoising”. Costs of Image Deblurring: On average, this tool takes about 667.42 milliseconds to run. It consumes approximately 444.91 MB of CPU memory and 3498.11 MB of GPU memory.
(Response Format) Provide a response in the similar format according to the following example: “Tool1, Tool2, Tool3”.
Non-Sequential Planning (Command) You need to act as a policy model that, given a task and a set of tools, determines the execution plan of tools that can be executed in sequential or parallel to solve the task. Your goal is to generate the tool plans that can optimize the task performance while minimizing the execution costs.
(Tool Information) Same as sequential planning.
(Response Format) Provide a response in the similar format according to the following example: “[‘Tool1’, [‘Task’], ‘Tool2’, [‘Tool1’], ‘Tool3’, [’Tool1’]]”. The meaning of this format is that the input data of Tool2 comes from the outputs of Tool1. Besides, “[‘Task’]” means that Tool1 depends on the input data provided by the task. Please generate plans strictly according to this format.
Table C.1: The prompts for zero-shot learning.
Scenarios Prompts
Sequential Planning (Command) Same as zero-shot learning.
(Tool Information) Same as zero-shot learning.
(In-Context Examples) Task: Given a low-resolution, noisy, blurry gray image, how to return the regular image step by step? Plan: Image Super Resolution, Image Denoising, Image Deblurring, Colorization
(Response Format) Same as zero-shot learning.
Non-Sequential Planning (Command) Same as zero-shot learning.
(Tool Information) Same as zero-shot learning.
(In-Context Examples) Task: Given a low-resolution grayscale image, how to (1) return the regular image, and (2) return the class label of the image in English? Plan: [‘Colorization’, [‘Task’], ‘Image Super Resolution’, [‘Colorization’], ‘Image Classification’, [‘Colorization’]]
(Response Format) Same as zero-shot learning.
Table C.2: The prompts for few-shot learning.

### v3

@@ 5.3 Ablation Study
Importance of different techniques in CAORL. The proposed fine-tuning algorithm CAORL plays an significant role in enabling CATP-LLM for cost-aware tool planning. Hence, we explore the contributions of various important techniques in CAORL to CATP-LLM, including context augmentation, intermediate feedback signals, and adaptive masking. As shown in Table 4, the context augmentation and intermediate feedback are essential for reducing execution costs, while adaptive masking ensures plan validity.
@@
Impacts of LLM parameter sizes. Finally, we explore the impacts of LLM sizes on CATP-LLM’s efficacy. We use TinyLlama-1B [61], Qwen-3B [22], Llama2-7B [49] and Qwen-14B [22] as backbone LLMs. As shown in Figure 7, Qwen-3B demonstrates comparable performance to Llama2-7B in sequential planning, but Llama2-7B significantly outperforms others in complex non-sequential tasks. This suggests that for relatively simple sequential tasks, smaller LLMs are sufficient for planning, while larger LLMs are recommended for more complex tasks.
@@ 6 Conclusion
In this paper, we propose the first cost-aware tool planning framework CATP-LLM. It features a tool planning language to enhance the LLM to create non-sequential plans of multiple branches for efficient concurrent execution. Building on this language, it further introduces a cost-aware offline reinforcement learning algorithm to fine-tune the LLM to optimize the trade-off between plan performance and costs. We then present OpenCATP, the first dataset for cost-aware planning evaluation. Experiments on OpenCATP show that CATP-LLM significantly outperforms GPT series in cost-aware tool planning, even with a 7B LLM as its backbone.
Acknowledgements
We thank the anonymous ICCV reviewers for their invaluable feedback. Duo Wu would also like to thank his colleagues in the lab, Shuzhao Xie, Rongwei Lu, Jiajun Luo and Xiaojie Li, for their constructive comments on improving the quality of this paper. This work was supported in part by National Key Research and Development Project of China (Grant No. 2023YFF0905502), National Natural Science Foundation of China (Grant No. 92467204, 62472249 and 62402264), and Shenzhen Science and Technology Program (Grant No. JCYJ20220818101014030 and KJZD20240903102300001).
@@ A Details of CAORL
Prompts of context augmentation. The complete prompt of context augmentation scheme is illustrated belows.
# Instructions
Your role is to serve as a policy model for tool planning. Given a task and a set of tools, you need to select some tools and determine the execution plan of tools that can be executed in sequential or parallel to solve the task. Your goal is to generate the tool plans that can optimize the task performance while minimizing the execution costs.
# Tool Information
Each tool has its own functionality. Executing a tool will incur some execution costs. Besides, the costs of each tool may vary based on the size of inputs. The following are the embedding features of each tool:
[Tool 1 Embeddings], ..., [Tool N Embeddings]
# Tool Cost Information
The cost features of each tool are as follows:
[Tool 1 Cost Embeddings], ..., [Tool N Cost Embeddings]
# Task Specifications and Input Attributes
Next, you will receive information about the task and the attributes of the current inputs.
Task specifications: Given an low-resolution blurry image, how to return a regular image?
Task input Attributes: {’has_image’: true, image_size’: (520, 780), has_text’: false, text_length’: None}.
Now please generate a tool plan.
Learning algorithms. The details of the decision transformer [8] based learning algorithm in CAORL are described as follows. Formally, the LLM takes the historical return, state, and action sequences to predict the next action:
@@ B Details of OpenCATP
B.1 Tool Set
Following OpenAGI [15], the tools OpenCATP for task solving consist of 10 open-source domain expert models of different functions. These models include: sentiment analysis [3], text summarization [27], machine translation [41], image classification [10], object detection [6], image colorization [62], image super-resolution [9], image denoising [59], image deblurring [59] and image captioning [11].
B.2 Evaluation Tasks
As mentioned in §$\S$4, OpenCATP includes 87 sequential tasks and 24 non-sequential tasks, each with 100 data samples. The sequential tasks in OpenCATP are constructed from OpenAGI [15], one of the most popular LLM sequential tool planning datasets. As for non-sequential tasks, we use the back-instruct method [46] for task construction. Specifically, we construct diverse non-sequential tasks by following steps: (1) create a tool graph based on the toolkit and dependencies between tools; (2) extract a sub-graph representing a non-sequential plan; (3) prompt the GPT-4 to generate the descriptions of a task that can be solved by this plan. Table B.1 provides several concrete examples of various types of tasks in OpenCATP.
Task Input Sample Ground Truth Example Plan
@@ B.2 Evaluation Tasks
Image-in-Image-out Task: Given a noisy blurry grayscale image, how to return the regular image step by step? [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Image-in-Text-out Task: Given a noisy blurry grayscale image, how to return the caption in English step by step? [Uncaptioned image] A woman stands in the dining area at the table. [Uncaptioned image]
@@
Image-in-Text-out Task: Given a blurry image, how to (1) return the label of the image in English, and (2) return the caption of the image in German step by step? [Uncaptioned image] (1) Braunbär; (2) Ein kräftiger Grizzly Bär ist im Hintergrund mit Gras zu sehen. [Uncaptioned image]
Image-in-Image-Text-out Task: Given a blurry noisy image, how to (1) return the regular image, and (2) return the label of the image in German step by step? [Uncaptioned image] (1) [Uncaptioned image]; (2) Schaf. [Uncaptioned image]
@@
B.3 Evaluation Metrics
Plan execution prices. OpenCATP utilizes execution prices to comprehensively reflect the overall costs of tool plans, inspired by the Function as a Service (FaaS) platforms. FaaS is a category of cloud computing that charges consumers in an event-driven manner, where consumers only need to pay for the time and resources used to run their functions, with no prices paid when the functions are idle [44]. This “pay-as-you-go” model can result in significant savings, especially for applications with intermittent usage or short-lived tasks, contributing to the market success of FaaS. According to recent statistics [23], the global market size of FaaS is estimated to be 17.70 billion USD in 2024 and will reach 44.71 billion USD by 2029.
@@ B.3 Evaluation Metrics
Here, $time$ denotes the execution time in milliseconds of tool $t$. $cpu_{cons}$ ($gpu_{cons}$) denotes the constant CPU (GPU) memory consumption in MB required to load the tool. $cpu_{inst}$ ($gpu_{inst}$) denotes the instant CPU (GPU) memory consumption in MB caused by the computation of the tool. $price\_cpu_{cons}$, $price\_cpu_{inst}$, $price\_gpu_{cons}$, $price\_gpu_{inst}$ calculate the prices in USD associated with the respective memory consumptions. $price\_per\_run$ denotes the prices in USD charged for each execution of the tool. Table B.2 lists the settings of the pricing parameters [^3] Note that the existing FaaS platforms do not provide pricing strategy for GPU resources. We then set the GPU prices as three times the CPU prices, according to the article in https://news.rice.edu/news/2021/rice-intel-optimize-ai-training-commodity-hardware.
@@
Plan execution time. For a sequential plan, the execution time is simply the sum of the execution time of each tool in the plan. For example, for plan ① in Figure 1, the execution time is $0.18+3.46+0.13=3.77$(s). For a non-sequential plan that can be executed in parallel, the execution time is defined as the maximum of the sum of tool execution time in each branch of the plan. For example, for plan ④ in Figure 1, the execution time is $\max(0.18+0.29+0.16,0.18+0.09)=\max(0.63,0.27)=0.63$ (s).
Task performance of plans. In OpenCATP, we mainly use ViT scores [15] and BERT scores [63] to calculate the performance scores of an executed tool plan on the target task. Specifically, ViT scores are used for tasks that involve image outputs, which measures the cosine similarity between the image generated by the plan and the ground-truth image. BERT scores are used for tasks that involve text outputs, which measures the cosine similarity between the generated texts and ground-truth texts.
@@ C Details of Experiments
C.1 Training and Testing Sets
For training sets, we select 2,760 and 480 samples from OpenCATP for sequential and non-sequential planning, respectively. As for testing sets, we select 720 and 480 samples from various tasks which are not present in the training sets for the evaluation of sequential and non-sequential planning, respectively.
C.2 Implementation of CATP-LLM
The details of implementing CATP-LLM on OpenCATP platform in the experiments are described as follows.
@@ C.2 Implementation of CATP-LLM
Details of TPL. For TPL, we assign each tool in OpenCATP with a unique tool token and a dependency token. The dependency token of a tool means accepting the outputs provided by this tool as inputs, as depicted in Figure 3. We also introduce a special dependency token $\langle task\rangle$ indicating accepting the task data as inputs. Note that in our experiments, we focus on resource dependencies, i.e., input-output dependencies. However, the concepts of TPL can be easily extended to other types of dependencies, such as order dependencies where there exist strict execution orders between tools.
@@
Details of CAORL. As for CAORL, we use the execution prices and task performance defined in §$\S$§B.3 to calculate the execution costs and plan performance in the reward function. Besides, to implement the context augmentation scheme, we need to derive the cost attributes of each tool. To achieve this, we categorize the input data in OpenCATP into $k=4$ size levels using k-means clustering, with the optimal number of clusters determined by the elbow method. We then profile the execution time and CPU/GPU memory consumption of each tool across varying input sizes, following the method outlined in §$\S$3.2.1. Based on these statistics, we calculate the overall costs for each tool using the execution prices defined in Eq.(B.3), which ultimately yield the tool cost attributes.
Details of fine-tuning. We apply the data generation method described in §$\S$3.2 to create 1,200 sequential tool plans and 780 non-sequential tool plans as training data. We then fine-tune CATP-LLM using Llama2-7B as the default backbone with LoRA rank 64. The total number of trainable parameters, including LoRA weights and learning tokens in TPL, are less than 1M, accounting for no more than 0.2% of the total parameters. Besides, it only takes CATP-LLM 9.5h and 6.2h to converge for 2 epochs in the sequential and non-sequential scenarios, when fine-tuned over a single 32GB V100 GPU. Hence, the fine-tuning overhead of CATP-LLM is small.
C.3 Implementation of Baselines
@@
Zero-shot learning. For zero-shot learning, we design two types of prompts for sequential planning and non-sequential planning, respectively. For sequential planning, we prompt the LLM to produce a tool sequence that can be executed sequentially to solve the target tasks. As for non-sequential planning, we instruct the LLM to generate a tool plan following the format similar to our TPL. This is because we find the LLM achieves poor performance in non-sequential planning without an appropriate format to generate non-sequential plans. Note that we also incorporate the average tool cost information into the prompts to instruct the LLM to create low-cost tool plans. The average tool cost information is derived from offline profiling results described in §$\S$§C.2. The prompts for zero-shot learning are shown below.
[Prompt for Sequential Planning]
# Instructions
You need to act as a policy model that, given a task and a set of tools, determines the sequence of tools that can be executed sequentially to solve the task. Your goal is to optimize the task performance while minimizing the execution costs.
# Tool Information
The information of each tool is provided as follows:
Object Detection: This tool identifies the names of objects in an image. It is generally used for object identification in the iamges. The input and output types of this tool are image and text, respectively. It can accept inputs from tools ’’Image Super Resolution’’, ’’Colorization’’’, ’’Image Deblurring’’, or ’’Image Denoising’’.
Costs of Object Detection: On average, this tool takes about 175.73 milliseconds to run. It consumes approximately 352.19 MB of CPU memory and 449.37 MB of GPU memory.
Image Deblurring: This tool can enhance the clarity of blurry images. It can be used for tasks that require improving image quality. Both the input and output types are images. It can accept inputs from tools ’’Image Super Resolution’’, ’’Colorization’’, or ’’Image Denoising’’.
Costs of Image Deblurring: On average, this tool takes about 667.42 milliseconds to run. It consumes approximately 444.91 MB of CPU memory and 3498.11 MB of GPU memory.
# Response Format
Provide a response in the similar format according to the following example: ’’Tool1, Tool2, Tool3’’.
[Prompt for Non-Sequential Planning]
# Instructions
You need to act as a policy model that, given a task and a set of tools, determines the execution plan of tools that can be executed in sequential or parallel to solve the task. Your goal is to generate the tool plans that can optimize the task performance while minimizing the execution costs.
# Tool Information
[Same as Sequential Planning]
# Response Format
Provide a response in the similar format according to the following example: [’Tool1’, [’Task’], Tool2’, [’Tool1’], Tool3’, [’Tool1’]]. The meaning of this format is that the input data of Tool2 comes from the outputs of Tool1. Besides, [’Task’] means that Tool1 depends on the input data provided by the task. Please generate plans strictly according to this format.
@@
Few-shot learning. The few-shot learning share similar prompts with zero-shot learning, except that we handcraft several high-quality demonstrations as in-context examples to augment the LLM for plan generation. The prompts for few-shot learning are shown below.
[Prompt for Sequential Planning]
# Instructions
[Same as Zero-Shot]
# Tool Information
[Same as Zero-Shot]
# In-Context Example
Task: Given a low-resolution, noisy, blurry gray image, how to return the regular image step by step?
Plan: Image Super Resolution, Image Denoising, Image Deblurring, Colorization
# Response Format
[Same as Zero-Shot]
[Prompt for Non-Sequential Planning]
# Instructions
[Same as Zero-Shot]
# Tool Information
[Same as Zero-Shot]
# In-Context Example
Task: Given a low-resolution grayscale image, how to (1) return the regular image, and (2) return the class label of the image in English?
Plan: [’Colorization’, [’Task’], Image Super Resolution’, [’Colorization’], Image Classification’, [’Colorization’]]
# Response Format
[Same as Zero-Shot]
@@
HYDRA. We reuse the pipeline and the prompt of HYDRA[24] in our experiments. To ensure fairness, we restrict the tool set of HYDRA to those available in OpenCATP. What’s more, similar to HuggingGPT, we have also added the tool cost information in the prompt of HYDRA.
@@
Note that the engines of GPT-3.5 and GPT-4 used for all prompt-based methods are gpt-3.5-turbo-0125 and gpt-4-turbo, respectively.
D Evaluate CATP-LLM on General Benchmark
Method Tool Planner Mean Accuracy (%) Average Number of Tools
CoT GPT-3.5 / 78.31 /
CoT GPT-4 / 83.99 /
Chameleon [34] GPT-4 86.54 3.40
Published Results (Above) $\blacktriangle$
CATP-LLM Llama2-7B 85.86 2.41
Table D.1: Comparing CATP-LLM with other methods on ScienceQA. CoT GPT-3.5/4 answer questions without using tools.
To validate the effectiveness of CATP-LLM on other benchmark except OpenCATP, we implement CATP-LLM on ScienceQA [34], a widely adopted multimodal benchmark for multi-choice question answering. We reused the tools in ScienceQA developed by Chameleon [35], a few-shot learning-based tool planning method. As ScienceQA lacks implementation for cost measurement (e.g., memory consumption), we use the number of tools in each plan as the indicator for tool planning costs. As shown in Table D.1, CATP-LLM outperforms CoT GPT-3.5 and GPT-4. Besides, CATP-LLM uses significantly fewer tools than Chameleon GPT-4 (a reduction of 29.11%) while its accuracy is only marginally lower than that of Chameleon GPT-4 (a decline of 0.68%). This highlights CATP-LLM’s advantage in cost-aware tool planning to substantially reduce tool costs without sacrificing performance. This finding on ScienceQA is consistent with that on OpenCATP, which demonstrates the generalization ability of CATP-LLM.
E Evaluate CATP-LLM on Various Hardware Equipment
Devices Methods Task Scores $\uparrow$ Exec. Price (\$) $\downarrow$ Runtime (s) $\downarrow$ QoP $\uparrow$
NVIDA RTX 3090 Sequential Planning
HuggingGPT (GPT-4) 0.665 0.097 1.108 0.246
Hydra (GPT-4) 0.603 0.061 0.654 0.247
CATP-LLM (Llama2-7B) 0.652 0.072 0.748 0.262
Non-Sequential Planning
HuggingGPT (GPT-4) 0.419 0.159 2.099 0.069
Hydra (GPT-4) 0.332 0.068 1.180 0.106
CATP-LLM (Llama2-7B) 0.566 0.138 0.761 0.161
NVIDA RTX 4090 Sequential Planning
HuggingGPT (GPT-4) 0.667 0.087 1.228 0.257
Hydra (GPT-4) 0.603 0.039 0.432 0.267
CATP-LLM (Llama2-7B) 0.651 0.045 0.487 0.286
Non-Sequential Planning
HuggingGPT (GPT-4) 0.419 0.103 1.413 0.118
Hydra (GPT-4) 0.332 0.049 0.844 0.122
CATP-LLM (Llama2-7B) 0.566 0.077 0.463 0.215
Table E.1: Evaluation on real-world commercial GPUs. Arrow $\uparrow$/$\downarrow$ means higher/lower is better.
According to the cost definition in Equation (B.1), the cost measurement (e.g., runtime and memory consumption) can be affected by hardware equipment. To validate the applicability of CATP-LLM across diverse hardware configurations, we test the plans of CATP-LLM and top baselines on various real-world commercial GPUs. We report the performance of these methods in Table E.1. We can see that CATP-LLM strikes the better balance between task scores and execution costs, leading to the highest QoP. This demonstrates the applicability of CATP-LLM in various hardware configurations.